OpenStudy (anonymous):

Will give medal? At the movie theatre, child admission is $5.90 and adult admission is $9.90 . On Saturday, twice as many adult tickets as child tickets were sold, for a total sales of $771.00 . How many child tickets were sold that day?

OpenStudy (anonymous):

x child tickets 2 x adult tickets 5.90 x + 9.90 ( 2x ) = 771 (5.90 + 19.80)x = 771 25.7 x = 771 x = 30

OpenStudy (anonymous):

Thank you how about? A Web music store offers two versions of a popular song. The size of the standard version is 2.4 megabytes (MB). The size of the high-quality version is 4.3 MB. Yesterday, the high-quality version was downloaded four times as often as the standard version. The total size downloaded for the two versions was 5684 MB. How many downloads of the standard version were there?